OpenStudy (anonymous):

calculate the determinant for 1² 2² 3² 4² 2² 3² 4² 5² 3² 4² 5² 6² 4² 5² 6² 7²

OpenStudy (anonymous):

Just to clarify is that indeed a matrix? :-)

OpenStudy (anonymous):

yes it is matrix

OpenStudy (anonymous):

actually i reduced this by eliminating the fisrt elements of first rows and reduced the matrix to 1 4 9 16 0 7 20 39 0 20 56 95 0 39 108 207

OpenStudy (anonymous):

\[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right] \rightarrow determinant=a*d-b*c\] \[\left[\begin{matrix}a & b & c \\ d & e & f \\ g & h & i\end{matrix}\right] \rightarrow determinate=a(ei-hf)-d(bi-hc)+g(bf-ec)\] \[\left[\begin{matrix}a & b & c & d \\ e & f & g & h \\ i & j & k & l\end{matrix}\right] \rightarrow determinate=???\] ² 2² 3² z 2\(^2\)